An ice skater rotating in place on (essentially fr
Answer #1

A)When a figure skater extends her arms, she increases the distance between the axis of rotation and some of her mass, increasing her moment of inertia. Since angular momentum is conserved, her rotational velocity must decrease to compensate. so kinetic energy decreases.

B)L = I* w ==> and L is constant.. so change in I = - (change in w)

also d(KE) = d(I) + 2*d(w) = d(I) - 2*d(I) = -d(I)

so change in kinetic energy = -(change in inertia)

the moment of inertia increases by 10%